A firm uses the weighted average method for process costing. It has 563 units in beginning WIP inventory at 31% complete with respect to conversion costs. It starts 25,827 units during the period and has 614 units in ending inventory (these units are 46% complete with respect to conversion costs). All direct materials are added at the beginning of the process.

What are total equivalent units with respect to direct materials?
Answer #1

Unit transferred out = 563+25827-614 = 25776

Equivalent unit of material = Unit transferred out+(Ending WIP*percent completion)

= 25776+(614*100%)

Equivalent unit of material = 26390